Listed below are the numbers of words spoken in a day by each member of eight different randomly selected couples. Complete pListed below are the numbers of words spoken in a day by each member of eight different randomly selected couples. Complete parts​ (a) and​ (b) below. Male 15 comma 951 27 comma 184 1425 7559 19 comma 399 15 comma 361 14 comma 070 26 comma 661 Female 24 comma 435 13 comma 458 17 comma 832 17 comma 567 13 comma 423 17 comma 199 16 comma 572 18 comma 095 a. Use a 0.05 significance level to test the claim that among​ couples, males speak fewer words in a day than females. In this​ example, mu Subscript d is the mean value of the differences d for the population of all pairs of​ data, where each individual difference d is defined as the words spoken by the male minus words spoken by the female. What are the null and alternative hypotheses for the hypothesis​ test?
Answer #1

Hypothesis:

Ho: \mu _{D} = 0

Ha: \mu _{D} < 0 .......males speak fewer words in a day than females.

Computational Table:

Male Female d = Male - Female d^2
15951 24435 -8484 71978256
27184 13458 13726 188403076
1425 17832 -16407 269189649
7559 17567 -10008 100160064
19399 13423 5976 35712576
15361 17199 -1838 3378244
14070 16572 -2502 6260004
26661 18095 8566 73376356
Total -10971.00 748458225

Calculation:

Di -10971.00 1371.38

νΣβ-Lie-V/748458225ーし 10235.88 8-1 rl

10235.85 3618.93

Test statistic:

d-11 D -1371.38-0 0.38 Sp 3618.93

Degrees of Freedom = n-1 = 8-1 = 7

Critical value:

ta,df = t0.05,8-1 1.8946

Conclusion:

Test statistic > Critical value, -0.38 > -1.8946, Therefore There is Fail to Reject Ho at 5% level of significance.

Therefore, There is Not sufficient evidence to males speak fewer words in a day than females.